the least value of cos^2 theta – 6sin theta + 3sin^2 theta + 2 is

Grade:11

1 Answers

Lab Bhattacharjee
8 years ago
\text{Using } \cos^2A=1-\sin^2A, \\ \text{the expression reduces to } 3(\sin\theta-1)^2